OpenStudy (anonymous):

The perimeter of a rectangle equalsteh sum of the lengths of its four sides. If the width of the rectangle is 9 inches, and its perimeter is 48 inches., find the length of the rectangle

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

p e n i s...

OpenStudy (mathstudent55):

In a rectanglew there are two equal lengths and two equal widths. P = L + L + W + W, where P = perimeter, L = Length, W = width, so P = 2L + 2W You are given P = 48 in. and W = 9 in., so substitute these values into the perimeter equation and solve for L: P = 2L + 2W 48 = 2L + 2(9) 48 = 2L + 18 30 = 2L 2L = 30 L = 15

OpenStudy (shane_b):

Your answer is more detailed so I'll just delete mine :)